OpenStudy (anonymous):

(2^8 ⋅ 3^−5 ⋅ 6^0)^−2 ⋅ 3 to the power of negative 2 over 2 to the power of 3, whole to the power of 4 ⋅ 2^28 (5 points)

OpenStudy (anonymous):

(2^8 x 3^5 x 6^0)^-2 x 3^-2/2^3 x 2^28

OpenStudy (igreen):

Is this it? @cami1231 \(\dfrac{(2^8 ⋅ 3^5 ⋅ 6^0)^{-2} ⋅ 3^{-2}}{2^3 ⋅ 2^{28}}\)

OpenStudy (anonymous):

\[(2^{8}*3^{5}*6^{0})^{-2} *3^{-2}/2^{3}*2^{28}\]

OpenStudy (anonymous):

I think so! Thanks so much! @igreen @juniro

OpenStudy (anonymous):

i might have other questions tho !

OpenStudy (anonymous):

(256*243*1)^-2*(1/72)*2^28 =512/531441=.000963

OpenStudy (anonymous):

actually i am wrong i went of your written equation and you missed a negative on 3^5 so your final answer should be 3359232

OpenStudy (anonymous):

Thanks @juniro

OpenStudy (anonymous):

oy yeah its 3^-5

OpenStudy (igreen):

\(\dfrac{(2^8 ⋅ 3^5 ⋅ 6^0)^{-2} ⋅ 3^{-2}}{2^3 ⋅ 2^{28}}\) Simplify the denominator by adding the exponents and keeping the base: \(\dfrac{(2^8 ⋅ 3^5 ⋅ 6^0)^{-2} ⋅ 3^{-2}}{2^{31}}\) Multiply -2 to all the exponents that are in the parenthesis: \(\dfrac{2^{-16} ⋅ 3^{-10} ⋅ 6^0 ⋅ 3^{-2}}{2^{31}}\) Simlify 3^-10 and 3^-2 by adding the exponents and keeping the base: \(\dfrac{2^{-16} ⋅ 3^{-12} ⋅ 6^0 }{2^{31}}\) Simplify 6^0: \(\dfrac{2^{-16} ⋅ 3^{-12}}{2^{31}}\) That's all you can simplify it without simplifying the exponents.

OpenStudy (anonymous):

Thank you ! @igreen but i made an error when typing. its supposed to be 3^-5

OpenStudy (igreen):

That's okay: \(\dfrac{2^{-16} ⋅ 3^{10} ⋅ 6^0 ⋅ 3^{-2}}{2^{31}}\) Simplify 6^0: \(\dfrac{2^{-16} ⋅ 3^{10} ⋅ 3^{-2}}{2^{31}}\) Simlify 3^10 and 3^-2 by adding the exponents and keeping the base: \(\dfrac{2^{-16} ⋅ 3^8}{2^{31}}\)

OpenStudy (anonymous):

thank you ! @igreen ♥

OpenStudy (igreen):

No problem!